Find the equation of the line joining the points
(─5, 2, 3) and (5, ─ 2, 3).

Answer by venugopalramana(3286)

Find the equation of the line joining the points
(─5, 2, 3) and (5, ─ 2, 3).
THE FORMULA IS
(X-X1)/(X2-X1)=(Y-Y1)/(Y2-Y1)=(Z-Z1)/(Z2-Z1)
(X+5)/(5+5)=(Y-2)/(-2-2)=(Z-3)/(3-3)
(X+5)/10=-(Y-2)/4=(Z-3)/0
OR
Z=3 AND
4(X+5)=-10(Y-2)..OR
2(X+5)+5(Y-2)=0...OR
2X+5Y=0